OpenStudy (anonymous):

find the quotient 4 6/7 divided by 5 1/4 enter as a mixed number and in simplest form

OpenStudy (anonymous):

now you did what we did in the previous problem and i will help you when you get stuck

OpenStudy (anonymous):

alrite

OpenStudy (anonymous):

20 168/28

OpenStudy (anonymous):

no, show me all of your work so i can tell you what mistake you did

OpenStudy (anonymous):

7*4 28 = 4 6/28 * 5 1/28= 20 168/28

OpenStudy (anonymous):

4 6/7=?

OpenStudy (anonymous):

i dont understand what your asking

OpenStudy (anonymous):

4 6/7= (7*4+6)/7

OpenStudy (anonymous):

6

OpenStudy (anonymous):

\[\frac{ 4\frac{ 6 }{ 7 } }{ 5\frac{ 1 }{ 4 } }\] \[\frac{ \frac{ 34 }{ 7 } }{ \frac{ 21 }{ 4 } }\] \[ \frac{ 34 }{ 7 } *\frac{ 4 }{ 21 } \] \[ \frac{ 34*4 }{ 7*21 } \] \[ \frac{ 136 }{ 147 } \]

OpenStudy (anonymous):

is this the answer

OpenStudy (phi):

this video shows how to change mixed numbers to improper fractions http://www.khanacademy.org/math/arithmetic/fractions/v/changing-a-mixed-number-to-an-improper-fraction

OpenStudy (phi):

once you learn how to change to improper fractions, you divide this video shows how http://www.khanacademy.org/math/arithmetic/fractions/v/dividing-fractions

OpenStudy (anonymous):

thank you